4x+y=-9 4x+2y=-10

Answer by checkley71(8403)

4X+Y=-9
-4X-2Y=10
-------------------------
-Y=1
Y=-1 ANSWER.
4X-1=-9
4X=-9+1
4X=-8
X=-8/4
X=-2 ANSWER.
4*-2+2*-1=-10
-8-2=-10
-10=-10